# Can slow pulsars in Milky Way globular clusters form via partial recycling?

### v1

@@ 1 Introduction
The vast majority of cluster pulsars have spin periods less than 303030\,30ms and low magnetic fields consistent with expectations for the standard recycling scenario (e.g., Phinney & Kulkarni, 1994). However, a subpopulation of slowly spinning pulsars with spin periods of $0.3-1\,$s and inferred magnetic fields of roughly $10^{11}-10^{12}\,$ has also been found (Lyne et al., 1996; Boyles et al., 2011). In the past year, two additional slow pulsars with spin periods of roughly $2$ and 444\,4s have been identified in M15 by the Five-hundred-meter Aperture Spherical Telescope (FAST; Zhou et al., 2024), bringing the current total up to six. With characteristic ages of $10^{8}\,$yr or less, these apparently young pulsars are impossible to explain via standard massive-star evolution, as the ≲$\lesssim 50\,$Myr lifetime of any plausible massive-star progenitor is significantly less than the typical 101010\,10Gyr age of their globular cluster hosts. Alternative formation channels have been proposed, in particular involving collapse of massive white dwarfs triggered by accretion from a binary companion (Tauris et al., 2013) or formed through a binary merger (Kremer et al., 2023a).
@@ 2 Comparing timescales for spin-up and dynamical encounters
(e.g., Phinney & Kulkarni, 1994), where $f\gtrsim 1$ is a factor that depends on the accretion details. We adopt the widely-accepted idea that the magnetic field decays rapidly in the early phases of the accretion process (Taam & van den Heuvel, 1986; Shibazaki et al., 1989; van den Heuvel & Bitzaraki, 1995; Tauris et al., 2012), so that the Alfvén radius is comparable to the neutron star radius throughout most of the spin-up process. Assuming a constant mass accretion rate $\dot{M}$ and $f\approx 1$, the time to spin up a neutron star to spin period $P$ is roughly
@@ 3 Further constraints on the minimum binary mass transfer rate
Observations of accreting X-ray pulsars provide further insight into the relevant minimum $\dot{M}$. The only known LMXBs (2A 1822-371, 4U 1626-67, GRO 1744-28, and Her X-1; Patruno & Watts, 2021) with measured spin periods above roughly 300300300\,300ms all have mass transfer rates of $10^{-10}-10^{-8}\,M_{\odot}$/yr (Chakrabarty et al., 1997; Rappaport & Joss, 1997; Coriat et al., 2012; Heinke et al., 2013; Bak Nielsen et al., 2017). The combination of high $\dot{M}$ and slow spins indicate these LMXBs have likely just started mass transfer. The absence of observed systems with both slow spins and low accretion rates implies systems that avoid $\dot{M}\gtrsim 10^{-10}\,M_{\odot}$/yr entirely likely do not form, consistent with expectations from evolutionary calculations discussed in the previous section.
@@ 5.1 Massive white dwarf mergers
Our favored model for forming young pulsars is collapse following the merger of a super-Chandrasekhar white dwarf binary, described in detail in Kremer et al. (2023a). The centers of old core-collapsed globular clusters are expected to harbor large populations of massive white dwarfs, as suggested by both simulations (e.g., Kremer et al., 2021) and observations (e.g., Vitral et al., 2022). Within these clusters, dynamical interactions form massive white dwarf binaries, and ultimately mergers, at a rate of roughly $10^{-7}$ mergers per year in the full Milky Way cluster population (Kremer et al., 2023a). More than $90\%$ of these white dwarf mergers have total mass in excess of the Chandrasekhar limit, which, along with their chemical compositions and mass ratios, make them strong candidates for collapse into neutron stars (King et al., 2001; Dessart et al., 2006; Schwab, 2021).
@@ 6 Conclusions and Discussion
K.K. acknowledges support from the Carnegie Theoretical Astrophysics Center (CTAC), where part of this work was conducted. C.S.Y. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) DIS-2022-568580. F.A.R. acknowledges support from NSF grant AST-2108624 and NASA ATP grant 1580NSSC22K0722.

### v3

@@ 1 Introduction
The vast majority of cluster pulsars have spin periods less than 303030\,30ms and low magnetic fields consistent with expectations for the standard recycling scenario (e.g., Phinney & Kulkarni, 1994). However, a subpopulation of slowly spinning pulsars with spin periods of $0.3-1\,$s and inferred magnetic fields of roughly $10^{11}-10^{12}\,$ has also been found (Lyne et al., 1996; Boyles et al., 2011). In the past year, two additional slow pulsars with spin periods of roughly $2$ and 444\,4s have been identified in M15 by the Five-hundred-meter Aperture Spherical Telescope (FAST; Zhou et al., 2024; Wu et al., 2024), bringing the current total up to six. With characteristic ages of $10^{8}\,$yr or less, these apparently young pulsars are impossible to explain via standard massive-star evolution, as the ≲$\lesssim 50\,$Myr lifetime of any plausible massive-star progenitor is significantly less than the typical 101010\,10Gyr age of their globular cluster hosts. Alternative formation channels have been proposed, in particular involving collapse of massive white dwarfs triggered by accretion from a binary companion (Tauris et al., 2013) or formed through a binary merger (Kremer et al., 2023a).
@@ 2 Comparing timescales for spin-up and dynamical encounters
(e.g., Phinney & Kulkarni, 1994), where $f\gtrsim 1$ is a factor that depends on the accretion details. We adopt the widely-accepted idea that the magnetic field decays rapidly in the early phases of the accretion process (Taam & van den Heuvel, 1986; Shibazaki et al., 1989; van den Heuvel & Bitzaraki, 1995; Tauris et al., 2012), so that the Alfvén radius is comparable to the neutron star radius throughout most of the spin-up process. Assuming a constant mass accretion rate $\dot{M}$ (already a conservative assumption; see Section 3) and $f\approx 1$, the time to spin up a neutron star to spin period $P$ is roughly
@@ 3 Further constraints on the minimum binary mass transfer rate
Observations of accreting X-ray pulsars provide further insight into the relevant minimum $\dot{M}$. The only known LMXBs (2A 1822-371, 4U 1626-67, GRO 1744-28, Her X-1, and GX 1+4; Patruno & Watts, 2021) with measured spin periods above roughly 300300300\,300ms all have mass transfer rates of $10^{-10}-10^{-8}\,M_{\odot}$/yr (Chakrabarty & Roche, 1997; Chakrabarty et al., 1997; Rappaport & Joss, 1997; Coriat et al., 2012; González-Galán et al., 2012; Heinke et al., 2013; Bak Nielsen et al., 2017). The combination of high $\dot{M}$ and slow spins indicate these LMXBs have likely just started mass transfer. The absence of observed systems with both slow spins and low accretion rates implies systems that avoid $\dot{M}\gtrsim 10^{-10}\,M_{\odot}$/yr entirely likely do not form, consistent with expectations from evolutionary calculations discussed in the previous section.
@@
As a final comment, we note that Equation (2) is derived assuming a steady, continuous accretion rate. However, accreting neutron stars observed with mean mass transfer rates ≲$\lesssim 10^{-10}\,M_{\odot}/$yr tend to accrete via episodic transient accretion (with much higher instantaneous accretion rates) owing to accretion disk instabilities. This can significantly reduce both $\Delta M$ and the timescale required to recycle neutron star spin (Bhattacharyya & Chakrabarty, 2017; D’Angelo, 2017; Kar et al., 2024), and thus $\tau_{P}$ in Equation (2) can be viewed as an upper limit on the true accretion timescale. More precise treatment of this episodic accretion does not change our basic conclusions – in fact, it makes the mismatch with the dynamical encounter timescale of Equation (3) even worse.
@@ 5.1 Massive white dwarf mergers
Our favored model for forming young pulsars is collapse following the merger of a super-Chandrasekhar white dwarf binary, described in detail in Kremer et al. (2023a). The centers of old core-collapsed globular clusters are expected to harbor large populations of massive white dwarfs, as suggested by both simulations (e.g., Kremer et al., 2021b) and observations (e.g., Vitral et al., 2022). Within these clusters, dynamical interactions form massive white dwarf binaries, and ultimately mergers, at a rate of roughly $10^{-7}$ mergers per year in the full Milky Way cluster population (Kremer et al., 2023a). More than $90\%$ of these white dwarf mergers have total mass in excess of the Chandrasekhar limit, which, along with their chemical compositions and mass ratios, make them strong candidates for collapse into neutron stars (King et al., 2001; Dessart et al., 2006; Schwab, 2021).
@@
Accretion induced collapse of massive white dwarfs via stable mass transfer from binary companions may provide a similar pathway for formation of young neutron stars in globular clusters (Tauris et al., 2013), and likely operates at similar rate to mergers (Kremer et al., 2021a). Although not in a globular cluster, the aforementioned 4U 1626-67 source (Giacconi et al., 1972) may exemplify this scenario. This system contains an apparently young ($P=7.66\,$s) accreting pulsar with a $10^{12}\,$ field in an ultracompact 42-minute binary with a ≲$\lesssim 0.026\,M_{\odot}$ white dwarf donor (Verbunt et al., 1990; Hemphill et al., 2021). However, given that all but one of the six slow pulsars known in globular clusters are single, the original binary companion would need to be disrupted by a subsequent dynamical encounter (Kremer et al., 2023a).
@@ 6 Conclusions and Discussion
Pulsars spun-up via accretion must occupy the space in the $P-\dot{P}$ diagram below the “spin-up line” indicating the shortest spin period that can be reached by accretion at the Eddington limit (Pringle & Rees, 1972). The fact that all six slow pulsars are found below or near the spin-up line (depending on assumptions about the uncertain accretion physics) has been touted as alternative evidence for the partial recycling scenario (Verbunt & Freire, 2014). If slow pulsars are born via white dwarf collapse, they need not obey the spin-up line limit. Future detection of a slow pulsar lying clearly above the spin-up line would provide strong evidence that neutron star formation is indeed ongoing in globular clusters.
@@
We thank the anonymous referee for their careful review and also thank Paulo Freire for comments on the manuscript. K.K. acknowledges support from the Carnegie Theoretical Astrophysics Center (CTAC), where part of this work was conducted. C.S.Y. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC) DIS-2022-568580. F.A.R. acknowledges support from NSF grant AST-2108624 and NASA ATP grant 1580NSSC22K0722.